OpenStudy (anonymous):

does anyone know how to put x^3-4x^2+2x+3 in quadratic form

OpenStudy (amistre64):

y = x-(4/3) ; x = y+(4/3) prolly not, but Decartes would do it

OpenStudy (agreene):

As I've mentioned, I basically know nothing about the quadratic form, but if i remember right--people like it more for 2nd degree polynomials, so... this factors to (-3+x) (-1-x+x^2) does that make it any easier?

OpenStudy (anonymous):

yea but how did u factor the -3+x out

OpenStudy (amistre64):

\[y^3+\frac{10}{3}y+\frac{25}{27}=0\] will tells us the roots for our sub; then we just -4/3 for it :)

OpenStudy (anonymous):

amistre64: how did you do this?

OpenStudy (amistre64):

i was reading in hawkings book today the chapter on descartes; he had a reduction method to get rid of the second term

OpenStudy (amistre64):

x = y + coeff/leading power

OpenStudy (amistre64):

it just shifts the graph really; then the roots of the new graph are just how far off you moved it

OpenStudy (amistre64):

5/3 - 4/3 = 1/3 as an original root

OpenStudy (anonymous):

|its difficult understanding that the degree is 4 but theres only 3 zeros because the original equation was \[x^4-7x^3+14x^2-3x-9 \] so what i did was find p/q first so the only rational zero was 3 bc i used synthetic division and i was left with x^3-4x^2+2x+3 so now im suppose to use the quadratic equation but i dont know how to factor something out in order to make it a quadratic

OpenStudy (amistre64):

lol, lets try it on that one :)

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=y%3D%28x%2B7%2F4%29%5E4%E2%88%927%28x%2B7%2F4%29%5E3%2B14%28x%2B7%2F4%29%5E2%E2%88%923%28x%2B7%2F4%29%E2%88%929 it got rid of the second term alright; now if only I knew what to do with it after that lol

OpenStudy (anonymous):

like i know how to do it but i dont remember how to factor it out to leave it as a quadratic can u help me on that

OpenStudy (amistre64):

you pretty much set up a pool of options as you best guesses

OpenStudy (amistre64):

\[\pm\frac{factor.last.term}{factor.first.term}\]

OpenStudy (amistre64):

since the leading term is a 1 that means that all our rational option are factors of 9 +- 1,3,9

OpenStudy (anonymous):

okay

OpenStudy (amistre64):

try out each term to see what zeros out x4−7x3+14x2−3x−9 0 1 -6 8 ------------------ 1 )1 -6 8 5 ... not it x4−7x3+14x2−3x−9 0 3 -12 6 9 --------------------- 3 )1 -4 2 3 0 ... 3 IS a root , and the remaining digits are our coeffs (x-3)(x^3 -4x^2 +2x +3) then we try the same technique for this new cubic

OpenStudy (anonymous):

okay im following

OpenStudy (amistre64):

+- 1,3 x^3 -4x^2 +2x +3 0 3 -3 -3 -------------------- 3) 1 -1 -1 0 .... that worked out nice, we got another zero (x-3)(x-3)(x^2 -x -1) the remaining is a quadratic that can be done up with whatever manner you desire

OpenStudy (amistre64):

synthetic division as it is called tends to be a bit smoother than longhand for this

OpenStudy (anonymous):

oh okay so u got 3 twice

OpenStudy (amistre64):

yes, which tells me that it touches the xaxis at x=3 and bend back again instead of actually going thru it

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=+x4%E2%88%927x3%2B14x2%E2%88%923x%E2%88%929 the graph shows that :)

OpenStudy (anonymous):

oh okay thanks dude i understand now u really helped dude